OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 16x - 2y = 74 2x - 2y = 4 A. (6, 3) B. (5, 0) C. (6, 0) D. (5, 3

OpenStudy (godlovesme):

solve for either y or x using either eqns

OpenStudy (zale101):

have you tried doing what Godlovesme mentioned? :)

OpenStudy (zale101):

@Tallan

OpenStudy (anonymous):

bout to right now

OpenStudy (zale101):

Ok ^_^

OpenStudy (zale101):

Step 1: Pick either one of the equation and solve for y

OpenStudy (anonymous):

y=8?

OpenStudy (godlovesme):

how did u get that?

OpenStudy (zale101):

Tallan, show step and mentioned what equation you used and the process of solving for y

OpenStudy (anonymous):

i plugged it into a calculator honestly. and it gave me 8 while using 16x-2y=74 @Zale101

OpenStudy (anonymous):

1) 16x - 2y = 74 2) 2x - 2y = 4 Rearrange both equations to isolate (-2y) then set these equal to each other: 74 - 16x = 4 - 2x 70 = 14x x = 5 Substitute into Eq(2): 2(5) - 2y = 4 10 - 2y = 4 6 = 2y y = 3 Check using Eq(1): 16(5) - 2(3) = 74 80 - 6 = 74 Answer (5,3)

OpenStudy (anonymous):

is that right on how to find it?

OpenStudy (zale101):

So you solved y for 2x - 2y = 4? \(\large 2x - 2y = 4\) rearrange the terms to isolate y, by subtracting both sides by 2x \(\large 2x \color{red}{-2x} - 2y = 4-\color{red}{2x}\) \(\large -2y=4-2x\) divide everything by -2 \(\large y=x-2\) How did you get y=8?

OpenStudy (anonymous):

i honestly don't know how i got 8

OpenStudy (anonymous):

i think i got confused on it.

OpenStudy (zale101):

I'm just confused, because i didn't know if you got y=8 by solving or plugging at the end ..

OpenStudy (zale101):

Ok, but you know how to do the substitution method? :)

OpenStudy (zale101):

You seem your're getting everything

OpenStudy (anonymous):

i think so. An if i have anymore trouble I will let you know.

OpenStudy (anonymous):

is that the right answer?

OpenStudy (zale101):

Yes

OpenStudy (godlovesme):

nice job guys @Zale101 & @Tallan

OpenStudy (godlovesme):

:)

OpenStudy (zale101):

Thanks :)

OpenStudy (anonymous):

Thanks :D :D

OpenStudy (godlovesme):

np :D

OpenStudy (zale101):

yw Tallan

OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Choose the correct ordered pair. 2x + 3y = 29 x = 4 A. (4, 7) B. (4, 11) C. (7, 4) D. (3, 2)

OpenStudy (anonymous):

could you maybe walk me through this just to make sure i got it @Zale101